Jake (80kg) and Mike (100kg) are standing 8 meters apart. What is the gravitational force between the two us us?


1
Expert's answer
2021-06-08T10:58:32-0400

"m_1= 80"

"m_2 =100"

"r = 8"

"G= 6.67*10^{-11}"

"\vec F=G\frac{m_1m_2}{r^2}"

"\vec F=6.67*10^{-11}\frac{80*100}{64}\approx8.4*10^{-9}"

"\text{Answer: }\vec F\approx8.4*10^{-9}\ N"